# STACKFEED: Structured Textual Actor-Critic Knowledge base editing with FEEDback

### v1

@@ Abstract
Large Language Models (LLMs) often generate incorrect or outdated information, especially in low-resource settings or when dealing with private data. To address this, Retrieval-Augmented Generation (RAG) uses external knowledge bases (KBs), but these can also suffer from inaccuracies. We introduce STACKFEED, a novel Structured Textual Actor-Critic Knowledge base editing with FEEDback approach that iteratively refines the KB based on expert feedback using a multi-actor, centralized critic reinforcement learning framework. Each document is assigned to an actor, modeled as a ReACT agent, which performs structured edits based on document-specific targeted instructions from a centralized critic. Experimental results show that STACKFEED significantly improves KB quality and RAG system performance, enhancing accuracy by up to 8% over baselines.
@@ 1 Introduction
Large Language Models (LLMs) often produce incorrect or outdated information, particularly in low-resource settings or when handling private data. Even if the information provided is accurate, LLMs can generate hallucinated or imaginary content alongside it (Maynez et al., 2020; Zhou et al., 2021). A promising solution to address these issues is the integration of retrieval components that extract relevant information from external knowledge sources, known as Retrieval-Augmented Generation (RAG) (Chen et al., 2017; Khandelwal et al., 2020; Guu et al., 2020; Izacard et al., 2022; Shi et al., 2023). For clarity, we will refer to these external knowledge sources as Knowledge Bases (KBs). However, KBs themselves can suffer from inaccuracies, incompleteness, or outdated content. To address these challenges, there is growing interest in Knowledge Editing (KE) techniques to enhance LLMs with up-to-date and accurate knowledge.
@@
Advancements in KE have focused on updating the model’s parameters (De Cao et al., 2021a; Meng et al., 2022; 2023), adding new parameters to model (Huang et al., 2023; Yu et al., 2024), and holding additional memory (Madaan et al., 2022; Wang et al., 2024a; b). Contrary to approaches that either update model parameters or add new parameters that require white-box access to LLMs, memory-based approaches can work with black-box access to LLMs. In similar line of thought, recently, KE approaches have also focused on refining the KBs themselves (Li et al., 2024). For example, the method proposed in Li et al. (2024) continuously updates KBs with new information, such as the current identity of the British Prime Minister. This approach demonstrates that directly editing the KB is more effective than simply adding new documents, which may coexist with outdated or inaccurate ones. Removing older documents is often not feasible, as only certain sections may be incorrect, while other parts could still provide valuable information for different queries. However, in applications like chatbots or code generation using API documentation, where updated information might not be readily available in document form, expert intervention can be crucial (Ramjee et al., 2024; Afzal et al., 2024). In such cases, expert feedback can be used to directly update the KB with accurate information when the LLM produces erroneous results.
  1. 1.
    Introduction of Feedback-Driven KB Editing: We present STACKFEED, a novel framework that refines the KB using structured edits based on expert feedback. This approach allows for direct, document-level updates without requiring access to LLM parameters, making it applicable to both white-box and black-box LLMs.
  2. 2.
    Multi-Actor, Centralized Critic Architecture: We design a multi-agent reinforcement learning framework where each actor is responsible for a specific document, and a centralized critic coordinates updates based on a global reward signal. This architecture ensures that document-level edits are consistent and contribute to the overall accuracy of the RAG system.
  3. 3.
    Parameterized Action Space for Document Editing: We propose a parameterized action space for each document-specific actor, enabling fine-grained control over edits, additions, and deletions within each document. This structured action space allows the actors to perform precise modifications based on expert feedback, resulting in a refined KB that better supports the RAG system.
  4. 4.
    Definition and Evaluation of KB Characteristics: We define desirable characteristics for KB refinement, including coherence, completeness, and generalizability, and introduce corresponding metrics to quantitatively assess these properties. These metrics provide a systematic way to measure the effectiveness of KB updates.
  5. 5.
    Empirical Evaluation and Performance Gains: We demonstrate that STACKFEED significantly improves the accuracy and reliability of the QA system in a variety of settings. Through extensive experiments, we show that incorporating expert feedback into document-level edits leads to a substantial reduction in error rates and enhances the KB’s ability to support accurate answer generation.
This paper is organized as follows: Section 2 reviews relevant prior work, while Section 3 presents an illustrative example to introduce and explain our approach. Section 4 details the proposed methodology, and Section 5 outlines the desired characteristics for the edited KB along with metrics for evaluation. Section 6 describes the experimental setup, and finally, Section 7 reports the results.
2 Related work
The STACKFEED framework addresses a key limitation of current RAG systems: the inability to dynamically update Knowledge Bases (KBs) without retraining or altering model parameters. Our work draws from research in Retrieval-Augmented Generation (RAG), Continual Learning, Model Editing, and feedback-driven prompt optimization, incorporating insights from Multi-Agent Reinforcement Learning (MARL) to propose an effective solution for KB editing.
Retrieval Augmented Generation (RAG): RAG systems enhance LMs by retrieving relevant knowledge from a KB based on the input query and appending it to the context, thereby addressing the limitations of standalone LMs that lack sufficient context and produce inaccurate answers (Chen et al., 2017; Khandelwal et al., 2020; Guu et al., 2020; Izacard et al., 2022; Shi et al., 2023). These systems dynamically construct contexts from unstructured KBs without modifying the LM’s internal parameters. STACKFEED further enhances RAG systems by refining the KB itself based on feedback, ensuring more accurate and up-to-date information.
Continual Learning: Continual Learning (CL) methods address the challenge of updating LMs in non-stationary environments by ensuring that new information is learned without forgetting previously acquired knowledge (Jin et al., 2022; Xu et al., 2023; Padmanabhan et al., 2023; Akyürek et al., 2024). These methods are often computationally intensive and require large-scale retraining, making them less suitable for scenarios requiring frequent updates or minimal computational resources. STACKFEED, by contrast, leverages expert feedback to perform direct edits to the KB, avoiding the need for extensive retraining.
Knowledge Editing: Knowledge Editing approaches fall into two categories: Model Editing, which modifies the LM parameters directly, and Input Editing, which updates the knowledge supplied to the model. While Model Editing efficiently alters specific facts using specialized secondary models or altering parameters (De Cao et al., 2021b; Meng et al., 2023), it struggles to ensure consistent updates across contexts (Onoe et al., 2023; Hua et al., 2024). In contrast, Input Editing modifies the KB itself, enabling updates to be reflected in outputs without changing model parameters (Madaan et al., 2022; Wang et al., 2024a; b; Li et al., 2024). STACKFEED builds on input editing techniques by leveraging expert feedback to refine the KB systematically, ensuring more accurate and consistent responses.
Prompt Optimization: With the advent of LMs, some recent works approximate gradients in text-based environments using LMs (Pryzant et al., 2023; Wang et al., 2023; Juneja et al., 2024; Gupta et al., 2024) for optimizing task prompts. STACKFEED is inspired by these approaches and generates textual reflections, similar to MetaReflection (Gupta et al., 2024) and Shinn et al. (2023), as proxies for gradients. It provides actionable guidance for document updates without the need for differentiable models. Additionally, STACKFEED adopts clustering strategies for feedback aggregation from works like UniPrompt (Juneja et al., 2024)- ensuring that actors receive coherent and non-redundant instructions.
Multi-Agent Reinforcement Learning (MARL): Multi-agent reinforcement learning (MARL) has been applied to various domains, with early research focusing on tabular methods (Busoniu et al., 2008; Canese et al., 2021; Gronauer & Diepold, 2022) and later expanding to deep learning techniques for high-dimensional inputs (Tampuu et al., 2017; Leibo et al., 2017). Studies have explored independent Q-learning (Tan, 1993), agent communication (Foerster et al., 2016; Das et al., 2017), and centralized training with decentralized execution (Gupta et al., 2017). However, most of these approaches do not address the critical challenge of multi-agent credit assignment. Actor-critic methods have been introduced to overcome this limitation by employing centralized critics with decentralized actors  (Foerster et al., 2018; Iqbal & Sha, 2019; Wang et al., 2021; Chen et al., 2023). STACKFEED extends such actor-critic framework to operate directly on textual content, using the centralized critic to decompose feedback into actionable textual gradients for each document-specific actor.
In the next section, we provide an example to illustrate the KB editing problem, while also providing an overview of STACKFEED.
3 Example and Overview
Figure 1 illustrates our technique applied to the ARKS Pony domain (Su et al., 2024a), where a knowledge base (KB) for the low-resource programming language Pony supports a natural language-to-code task. Due to Pony’s rarity, language models often generate code that fails to compile. To address this, we use the Pony compiler as an expert to provide feedback in the form of compile errors.
1
2
Centralized Feedback Analysis: We analyze compile errors to generate reflections that explain why the errors occurred. For instance, if the apply method in the Array class is partial and may raise an error, the reflection suggests adding a ? to handle potential failures. These reflections are matched to the documents they pertain to, refining the understanding of errors.
3
Distributing Gradients: Reflections are generalized into gradients, which summarize modifications needed for each document. For example, the theme might be the partial nature of functions like apply and update, which need better error handling in the documentation.
4
Generating Edit Actions: Gradients are converted into structured edit actions, such as adding or modifying content in specific sections of the documents.
5
Re-evaluation and MCTS Search: After edits are applied, the KB is re-evaluated, generating new feedback and a reward score. This score guides a Monte Carlo Tree Search (MCTS) to explore different states of the KB, iterating through steps 1- 4 to progressively refine the KB and improve the system’s overall performance.
@@ 3 Example and Overview
Feedback Reflection: The output code contains the following errors: 1. The array access calls nums(i) and nums(j) are not marked as partial, but they should be as the apply function in the Array class is partial and can throw errors. The corrected code is: nums(i)? = nums(j)? To further illustrate the fix, consider the following simplified example: Code with error: fun get(arr: Array[I32], index: USize): I32 ? => arr(index) Fixed code: fun get(arr: Array[I32], index: USize): I32 ? => arr(index)? ... Selection: builtin-array.md
Reasoning: This file contains information about the Array class, including apply...
Collected selection reasoning for builtin-array.md Reflection and Selection Document-wise Partial Gradient: builtin-array.md is a KB document detailing the use of the methods and functions of the Array class...
Method Availability and Error Handling: The current file lacks explicit guidance on the partiality of methods and the use of ? operator. The omission can lead developers to... The file should accurately reflect the partiality of methods like apply and update...
@@
When working with arrays in Pony, it is important to understand that some methods are partial, meaning the may not succeed in all cases and might raise an error. To safely handle these errors, Pony provides the ? operator...
The apply method on arrays is partial and will raise an error in the index is out of bounds...
@@
‘‘‘pony try my_array.apply(42) == 7 end ‘‘‘ In this snippet, if the index 42 is out of bounds the ‘apply‘ method will raise an error, which is then handled by the ‘try‘ block, preventing the program from crashing.
Generate Edits
@@
Figure 1: Example of the STACKFEED in the ARKS Pony scenario
@@ 4 Methodology
We will start by describing a typical Retrieval-Augmented Generation (RAG) system over unstructured Knowledge Bases.
Errors in such systems can arise from multiple components: 1) the LLM $\mathcal{B}$ might fail to reason correctly over the provided information, 2) the retriever $R$ might not select the right set of relevant documents from $\mathcal{K}$, or 3) the knowledge base $\mathcal{K}$ itself might contain incorrect or incomplete information. We assume an expert is monitoring the system, identifying when answers are incorrect, determining which component is at fault, and providing feedback on why the answer is incorrect and what the correct answer must be.
This work focuses on scenarios where incorrect answers result from issues in the Knowledge Base ($\mathcal{K}$). Our goal is to improve $\mathcal{K}$ by addressing mistakes in $\mathcal{K}$ and filling in missing information based on expert feedback, thus enhancing the RAG system’s performance on future queries.
4.1 Problem Formulation
We are provided with a training set $T=\{(q_{i},o_{i},c_{i},f_{i})\}_{i=1}^{l}$, where $q_{i}$ is a user query, $o_{i}$ is the RAG system’s answer, $c_{i}$ is the correct answer, and $f_{i}$ is an optional expert feedback on incorrect answers. We also assume access to a scoring function $g$, which compare $o_{i}$ and $c_{i}$ to output a score. The objective is to optimize the knowledge base $\mathcal{K}$ to maximize the sum of the scores for all queries in the training set:
$\mathcal{K}^{*}=\arg\max_{\mathcal{K}}\frac{1}{|T|}\sum_{(q_{i},a_{i},c_{i},f_{i})\in T}g(\mathcal{B}(q_{i},\Gamma(q_{i},\mathcal{K})),c_{i})$ (1)
In the next section, we show how such an objective can be seen as a state search problem.

### v2

@@ Abstract
Large Language Models (LLMs) often generate incorrect or outdated information, especially in low-resource settings or when dealing with private data. To address this, Retrieval-Augmented Generation (RAG) uses external knowledge bases (KBs), but these can also suffer from inaccuracies. We introduce STACKFEED, a novel Structured Textual Actor-Critic Knowledge base editing with FEEDback approach that iteratively refines the KB based on expert feedback using a multi-actor, centralized critic reinforcement learning framework. STACKFEED defines a ReACT actor agent on each document to perform structured edits based on document-specific targeted instructions. Experimental results showcase that STACKFEED significantly improves KB quality and performance of the RAG system. We evaluate STACKFEED on low-resource programming problems, modified python packaged and factual question-answering tasks.
@@ 1 Introduction
Large Language Models (LLMs) often produce incorrect or outdated information, particularly in low-resource settings or when handling private data. Even if the information provided is accurate, LLMs can generate hallucinated or imaginary content alongside it (Zhang et al., 2025; Maynez et al., 2020; Zhou et al., 2021). A promising solution to address these issues is the integration of retrieval components that extract relevant information from external knowledge sources, known as Retrieval-Augmented Generation (RAG) (Chen et al., 2017; Izacard et al., 2022; Shi et al., 2023; Li et al., 2025b). For clarity, we will refer to these external knowledge sources as Knowledge Bases (KBs). However, KBs themselves can suffer from inaccuracies, incompleteness, or outdated content. To address these challenges, there is growing interest in Knowledge Editing (KE) techniques to enhance LLMs with up-to-date and accurate knowledge.
@@
Advancements in KE have focused on updating the model’s parameters (De Cao et al., 2021a; Meng et al., 2022, 2023), adding new parameters to model (Huang et al., 2023; Yu et al., 2024), and holding additional memory (Madaan et al., 2022; Wang et al., 2024a, b). Contrary to approaches that update model parameters or add new parameters that require white-box access to LLMs, memory-based approaches can work with black-box access to LLMs. In a similar line of thought, recently, KE approaches have also focused on refining the KBs themselves (Li et al., 2025a). For example, the method proposed by Li et al. (2025a) continuously updates KBs with new information when presented with a document containing the exact information to be updated, such as updating the current identity of the British Prime Minister in the KB when the news of election results is provided. This approach demonstrates that directly editing the KB is more effective than simply adding new documents, which may coexist with outdated or inaccurate ones. Removing older documents is often non-trivial, as only certain sections may be incorrect, while other parts could still provide valuable information for different queries.
However, in applications such as chatbots or code generation tools that rely on API documentation, up-to-date information may not always be readily available in structured documents (Ramjee et al., 2024; Afzal et al., 2024; Liu et al., 2025b). In these scenarios, expert feedback becomes essential—not only for correcting erroneous outputs from the LLM but also for directly updating the underlying knowledge base (KB) with accurate information. This need is particularly pressing in live systems that depend on real-time, reliable data. Domains like healthcare (Ramjee et al., 2024), legal tech (Liu et al., 2025b), and financial services demand high precision and instant updates. Ensuring continuous and trustworthy KB revisions is therefore critical to maintaining the safety, effectiveness, and reliability of RAG applications in such high-stakes environments.
@@
Feedback Reflection: The output code contains the following errors: 1. The array access calls nums(i) and nums(j) are not marked as partial, but they should be as the apply function in the Array class is partial and can throw errors. The corrected code is: nums(i)? = nums(j)? To further illustrate the fix, consider the following simplified example: Code with error: fun get(arr: Array[I32], index: USize): I32 ? => arr(index)
Fixed code:
fun get(arr: Array[I32], index: USize): I32 ? => arr(index)?
Selection: builtin-array.md
Reasoning: This file contains information about the Array class, including apply…
Collected selection reasoning for builtin-array.mdCentralized Critic Document-wise Partial Gradient: builtin-array.md is a KB document detailing the use of the methods and functions of the Array class…
Method Availability and Error Handling: The current file lacks explicit guidance on the partiality of methods and the use of ? operator. The omission can lead developers to… The file should accurately reflect the partiality of methods like apply and update
@@
When working with arrays in Pony, it is important to understand that some methods are partial, meaning the may not succeed in all cases and might raise an error. To safely handle these errors, Pony provides the ? operator…
The apply method on arrays is partial and will raise an error in the index is out of bounds…
@@
“‘pony try my_array.apply(42)? == 7 end “‘ In this snippet, if the index 42 is out of bounds the ‘apply‘ method will raise an error, which is then handled by the ‘try‘ block, preventing the program from crashing.
Per-Document Actor
@@
Figure 1: Example of the STACKFEED pipeline in the ARKS Pony scenario. We explain the example in more detail in appendix A.3
  1. 1.
    Introduction of Feedback-Driven KB Editing: We present a novel framework that refines the knowledge base using structured edits based on oracle or expert feedback.
  2. 2.
    Definition and Evaluation of KB Characteristics: We define desirable characteristics for knowledge base refinement, including coherence, completeness and introduce corresponding metrics to quantitatively assess these properties.
  3. 3.
    Empirical Evaluation and Performance Gains: We demonstrate that STACKFEED significantly improves the accuracy and reliability of RAG system in a variety of settings.
2 Related work
The STACKFEED framework addresses a key limitation of current RAG systems: the inability to dynamically update Knowledge Bases (KBs) without retraining or altering model parameters. Our work draws from research in Retrieval-Augmented Generation (RAG), Continual Learning and incorporating insights from Multi-Agent Reinforcement Learning (MARL) to propose an effective solution for KB editing.
Retrieval Augmented Generation (RAG): RAG systems enhance LMs by retrieving relevant knowledge from a KB based on the input query and appending it to the context, thereby addressing the limitations of standalone LMs that lack sufficient context and produce inaccurate answers (Chen et al., 2017; Khandelwal et al., 2020; Guu et al., 2020; Izacard et al., 2022; Shi et al., 2023). These systems dynamically construct contexts from unstructured KBs without modifying the LM’s internal parameters. STACKFEED further enhances RAG systems by refining the KB itself based on feedback, ensuring more accurate and up-to-date information. Recent works showcase the failure of RAG due to inconsistencies in the retrieved documents Xiang et al. (2024); Wang et al. (2025).
Knowledge Editing: Knowledge Editing approaches fall into two categories (Yao et al., 2023b): Model Editing, which modifies the LM parameters directly, and Memory Based, which updates the knowledge supplied to the model. While Model Editing efficiently alters specific facts using specialized secondary models or altering parameters (De Cao et al., 2021b; Meng et al., 2023), it struggles to ensure consistent updates across contexts (Onoe et al., 2023; Hua et al., 2024). In contrast, Memory Based methods either add an auxiliary counterfactual model trained on new information (Mitchell et al., 2022) or they modify the KB itself, enabling updates to be reflected in outputs without changing model parameters (Wang et al., 2024b; Li et al., 2025a). STACKFEED builds on memory based techniques by leveraging expert feedback to refine the KB systematically, ensuring more accurate and consistent responses.
Prompt Optimization: With the advent of LMs, some recent works approximate gradients in text-based environments using LMs (Pryzant et al., 2023; Wang et al., 2023; Kirtania et al., 2024; Gupta et al., 2024) for optimizing task prompts. STACKFEED is inspired by these approaches and generates textual reflections, similar to MetaReflection (Gupta et al., 2024) and Shinn et al. (2023), as proxies for gradients. It provides actionable guidance for document updates without the need for differentiable models. Additionally, STACKFEED adopts clustering strategies for feedback aggregation from works like UniPrompt (Juneja et al., 2024)- ensuring that actors receive coherent and non-redundant instructions.
Figure 2: a) MCTS (Monte Carlo Tree Search) planning for state search. The tree structure enables strategic planning for STACKFEED. b) A simplified state transition example. Upon receiving a reward from the environment (or expert) on the given state of the knowledge base (KB) \$s_0\$⃝, a centralized critic ① generates a reflection on observed failures to calculate the textual gradient. The critic uses this reflection to select documents responsible for the error and ② assigns credit to actors in the form of document-wise reflections. The actors then iteratively edit the documents to reach state \$s_4\$⃝.
3 Problem Formulation
Typical RAG systems assume that the information present in the retrieved documents is correct and consistent. Our work focuses on scenarios where incorrect answers are generated due to issues in the retrieved documents from a Knowledge Base ($\mathcal{K}$).
More formally, we define the $\mathcal{K}$ as a collection of documents $D_{i}$ for $i=1,\ldots,N$. Each document $D_{i}$ can be represented as a set of chunks $c_{ij}$. The state of $\mathcal{K}$ is the specific configuration of all chunks within it.
For a given query $q$, such as the code generation Task in Figure 1, a retriever fetches a set of relevant documents $\Gamma(q,\mathcal{K})$. In the example, this corresponds to the Retrieved Documents list, which includes builtin-array.md and collections-persistent-vec.md.
An LLM, $M$, then generates a response $r$ based on the query and the retrieved documents, i.e., $r=M(q,\Gamma(q,\mathcal{K}))$. The initial Output Program in Figure 1 is an instance of such a response.
However, this response $r$ may be incorrect. We obtain feedback on the correctness of $r$, for instance, from Expert (Compiler) Feedback as shown in Figure 1. This feedback reveals flaws in the generated program that stem from deficiencies in $\mathcal{K}$. The expert is used as a scoring function, $g$ which evaluates whether a response $r$ is correct or not.
Our goal is to optimize the Knowledge Base state $\mathcal{K}$ to state $\mathcal{K}*$ which maximizes the scores for the responses for a batch of queries $\mathcal{Q}$ by learning from expert feedback:
$\mathcal{K}^{*}=\arg\max_{\mathcal{K}}\frac{1}{|\mathcal{Q}|}\sum_{q_{i}\in\mathcal{Q}}g(\mathcal{M}(q_{i},\Gamma(q_{i},\mathcal{K})))$ (1)
@@ 4 Methodology
We propose STACKFEED, an agent that employs Monte Carlo Tree Search (MCTS) to search for an optimal state of the knowledge base (KB). STACKFEED utilizes a multi-actor, centralized critic reinforcement learning architecture to guide transitions between KB states. This setup enables efficient exploration of a large, structured edit space Wang et al. (2023); Gupta et al. (2024), facilitating strategic and interpretable knowledge refinement.
The use of a multi-actor architecture with a centralized critic is central to STACKFEED’s design. In this formulation, each document within the KB is managed by a dedicated actor responsible for localized edits, while a centralized critic provides joint feedback by aggregating signals across all actor-document interactions. Our design aligns with recent findings from Lyu et al. (2024), who show that history–state centralized critics—critics conditioned on both joint observation histories and global state—can offer accurate and stable policy gradients, particularly in settings with partial observability and distributed decision-making.
In our context, where agents must coordinate to revise distinct portions of a shared KB based on sparse oracle feedback, this centralized view enables effective credit assignment. Specifically, we incorporate mechanisms inspired by counterfactual multi-agent policy gradients (COMA) (Foerster et al., 2018) to attribute responsibility for errors to specific documents and actors. The centralized critic leverages these credit signals to generate high-quality textual reflections that guide each actor’s editing trajectory. This design allows STACKFEED to maintain coherence across KB documents while iteratively improving task-specific correctness and completeness.
By combining decentralized editing with centralized, feedback-driven evaluation, STACKFEED learns to make targeted and interpretable updates to the KB. This architecture is particularly suited for real-world retrieval-augmented generation (RAG) settings, where maintaining consistency and relevance across independently edited documents is critical for downstream performance.